csc^2y tan^2y - 1 = tan^2y
Answer by mananth(16949)

csc^2y tan^2y - 1 = tan^2y

(1/sin^2y)(sin^2y)/cos^2y) -1
(1/cos^2y)-1

(1- cos^2y)/cos^2y but (sin^2y+cos^2=1)

= sin^2y/cos^2y
= tan^2y